OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

i need help i have no idea what to do

OpenStudy (anonymous):

Do you know what a quadratic function is?

OpenStudy (anonymous):

f(x)=a(x-h)^2+k

OpenStudy (anonymous):

Just give this a read, it will help you. http://www.mathsisfun.com/algebra/quadratic-equation-graphing.html

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I looked at the link you sent me and I get how to solve the equation but I'm not sure how they are graphing it.

OpenStudy (anonymous):

Well finish of the equation, and then post it here, I'll help you graph it then.

OpenStudy (anonymous):

ok so I got f(x)=a(x-h)^2+k My birthday is January 9th so I would use that and plug it into the equation. a can be anything between 4 and -4 so I will use 2. f(x)=2(x-1)+9 2x-2+9 2x+7

OpenStudy (anonymous):

To get it in standard form it has to look like this y=ax^2+bx+c

OpenStudy (anonymous):

So I would simplify it by getting rid of the parenthesis f(x)=2(x-1)+9 f(x) is the same thing as y so y=2x-2+9 y=2x+7

OpenStudy (anonymous):

so then would I just graph y=2x+7?

OpenStudy (anonymous):

Did you get the link I messaged you, because you're not doing it right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Im not sure what I did wrong fo rthe first question it is just asking for a formula not to solve it and it has to have h and k in it

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the second question is to take that formula and everything i have substituted into it and put it in standard form

OpenStudy (anonymous):

then the 3rd question I need help with the graphing

OpenStudy (anonymous):

\[f(x)=a(x-h)^2+k ~~~~ general~quadratic~ form\] Let a = 1 And fill in the rest

OpenStudy (anonymous):

1.Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2.Using complete sentences, explain how to convert your birthday function into standard form. 3.Graph your function. Include your graph below. 4.Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Lets do this in order so for 1 you just need to make the equation like I showed you above.

OpenStudy (anonymous):

Put h as your month and k as the day, so fill that in and you'll be done 1, once you do that, we'll move onto number 2.

OpenStudy (anonymous):

My birthday is january 9th f(x)=a(x-h)^2+k f(x)=1(x-1)^2+9

OpenStudy (anonymous):

Alright good, so that's all you have to do for #1 Lets move onto #2 now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This one you should be able to do on your own, all you have to do is take the form you have and make it into the standard form \[ax^2+bx+c = 0\]

OpenStudy (anonymous):

Yes this time you foil.

OpenStudy (anonymous):

f(x)=1(x-1)^2+9 1(x-1)(x-1)+9 something like this

OpenStudy (anonymous):

Yes, foil it out

OpenStudy (anonymous):

You can ignore the a = 1 part, we chose it because of this reason, it's easier for us :). So what's (x-1)(x-1)+9?

OpenStudy (anonymous):